I have tried this every way that I can think of, and I still cannot complete it. Please help!! The equation is as follows:
x+y/z

x= 14 2/5
y= 5
Z= 5 7/12


Answer by Alan3354(69443)

The equation is as follows:
x+y/z
x= 14 2/5
y= 5
Z= 5 7/12
-------------
It's not an equation, there's no equal sign. x + y/z is an expression, or a binomial.
I think you're asked to evaluate it for the values given.
answer = x + y/z
Use a, it saves typing
x = x + y/z
-------------------
Do the y/z first.
y/z = 5/(5 7/12) = 5/(67/12)
Invert the 67/12 and multiply.
y/z = 5*(12/67) = 60/67
--------------------
x = 14 2/5 = 72/5
a = 72/5 + 60/67
Find the common denominator by multiplying the 2 DENs. That ALWAYS works, tho in some cases it doesn't give the Least Common Denominator (LCD) that teachers fret about. If the CD is not the LCD, you just have to reduce the fraction. Don't waste time on get the LCD unless it's asked for.
----------
In this problem, it's 5*67 = 335
a = 72/5 + 60/67 = 72*67/DEN + 60*5/DEN
a = 4824/DEN + 300/DEN
a = 5124/335 That's it, unless you want to reduce it.
a = 15 99/335